A case study in the chapter analyzed purchasing-power parity for several countries using the price of Big Macs. Here are data for a few more countries: 


For each country, select the predicted exchange rate of the local currency per U.S. dollar. (Hint: Recall that the U.S. price of a Big Mac was $5.58.) 

  Country      Price of a Big Mac   Predicted Exchange Rate   Actual Exchange Rate

  Costa Rica      2,290 colones                              607 colones/$

  Pakistan         460 rupees                                 139 rupees/$

 Thailand           119 baht                                   32 baht/$

  China            20.9 yuan                                  6.85 yuan/$

 New Zealand       6.20 NZ$                                   1.48 NZ$/$


 According to purchasing-power parity, the predicted exchange rate between the Pakistani rupee and the New Zealand dollar is _______ rupees per New Zealand dollar. However, the actual exchange rate is _______  rupees per New Zealand dollar.
Answer #1

(1) Predicted exchange rate:

Costa Rica: 2,290 / 5.58 = 410.39 colones/$

Pakistan: 460 / 5.58 = 82.44 rupees/$

Thailand: 119 / 5.58 = 21.33 baht/$

China: 20.9 / 5.58 = 3.75 yuan/$

New Zealand: 6.2 / 5.58 = 1.11 NZ$/$

(2) Predicted exchange rate is 74.27 rupees per NZ dollars (= 82.44 / 1.11). Actual exchange rate is 93.92 rupees per NZ dollars (= 139 / 1.48).